OpenStudy (anonymous):

Let P(C) = 0.4 Let P(D) = 0.5 Let P(C|D) = 0.6 a. Find P(C AND D) . d. Find P(C OR D) . e. Find P(D|C).

OpenStudy (anonymous):

\[P(C/D)=\frac{ P \left( C and D \right) }{P \left( D \right) }\] \[0.6=\frac{ P \left( A and D \right) }{ 0.5 },\] P (C and D)=0.6*0.5= 0.3 \[P \left( C or D \right)=P \left( C \right)+P \left( D \right)-P \left( C and D \right)\] \[P \left( \frac{ D }{ C } \right)=\frac{ P \left( C and D \right) }{P \left( C \right) }\] CALCULATE.